x + 1/x – 2 = 0
x = 

Grade:9

4 Answers

noogler
8 years ago
-1
A M S ARUN KRISHNA
8 years ago
x + 1/x + 2 = 0
(x2 + 1 + 2x) /x = 0
x2 + 1 + 2x = 0
(x + 1)2 = 0 
x + 1 = 0
x = – 1
noogler
8 years ago
x+1=0(x-2)...........x=-1
 
it is easy.i think u should try these kind of sums urself :)
ashish
8 years ago
x+1/x=2
S O B S
x.x+1/x.1/x=2
by equating both equations
x=i